# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

### v2

@@ 1. Introduction
When we iterate a homomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
@@
There is yet another classification of these points based on the nature of their orbits; viz., the Filled in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. $K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]). The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain. Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions. We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is also transcendental.
@@
Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$.
The following results are proven in this paper:
@@
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed functions is contained in the union of the bungee set of functions used for composition. The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the union of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cup RP(g)$.
@@
We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18]. Also, $\mathcal{S}\subset\mathcal{B}$. $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition arewith $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, for such a case, all sets of composed functions are equal to the functions taken for composition.
@@ 1.1. Observations
    Both sets of composite function is contained in union of individual function (to be proved in this paper).
@@ 2. Main results
To begin with the commuting entire functions, we first understand the nature of these sets when functions have no asymptotic values in their singular set.
@@
But assuming $g(z)$ does not have any finite asymptotic value, so $|g(f^{n}(z_{0})|\rightarrow\infty$.
Since $f\,\&\,g$ are commuting so, $|g(f^{n}(z_{0})|$ = $|f^{n}(g(z_{0})|\rightarrow\infty.$ This implies that $g(z)$ carries $I(f)$ inside itself or in other words, $I(f)$ is forward invariant under $g$. It is easy to see that $g^{-1}(I(f))\subset I(f)$. Hence, $I(f)$ is completely invariant under $g$. As $f\,\&\,g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well. Hence, $I(f)$ and $I(g)$ is completely invariant under $f\circ g$ as well. ∎
@@
Suppose $f$ and $g$ are commuting entire functions having no asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$.
@@
$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|f^{m_{k}}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, thereby proving the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implis that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values.
Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know that $I(f)$ and $K(f)$ are completely invariant under $g$ (2.1) and hence, $BU(f)$ is also completely invariant under $g$. ∎
@@
The filled-in Julia set of the composite function $f\circ g$ cannot be larger than intersection of filled in Julia set of the individual functions having no finite asymptotic values.
@@
We assume that $K(f\circ g)\supset K(f)\cap K(g)$. Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$. If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, these points cannot be bounded as one function carries the points of other functions to itself, meaning $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory like bungee points nor will tend to $\infty$. Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$. However, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$, which cannot be bounded ($f(BU(g)\subset BU(g)$ and $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded. As these functions have no finite asymptotic value, so $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ will be unbounded in the case when $w_{0}\in I(g).$) Similarly, such arguments may be given when $w_{1}\in B$. Thus, we arrive at a contradiction. Hence, $K(f\circ g)\subset K(f)\cap K(g)$. ∎
@@
Let $z_{0}\in BU(f\circ g)$ where $f\ \&\ g$ do not have any finite asymptotic value. To begin with let there be two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ such that $|(f\circ g)^{m_{k}}(z_{0})|<R$ for some $R$ and $|(f\circ g)^{n_{k}}(z_{0})|\rightarrow\infty.$
So, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$. Similarly for $f,$ since they are commuting so, $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$.
Now, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ exhibits bungee behaviour. If we have the second case, it means $z_{0}\in K(g)$ (This will imply all subsequences of ($g^{j}(z_{0})$) is bounded where $j$ is an arbitrary index). To understand this, we take $w_{0}\in K(g)$. We observe $g^{j}\circ f^{2j}\circ g^{j}(w_{0})\in g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded for $j\rightarrow\infty$. Hence, this case isn’t possible. Now, when we have the third and the first case, then both essentially states that $z_{0}\in BU(g)$. From (2.2), we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected. And due to commutation, we also conclude that $z_{0}\in BU(f)$ as well. Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
Given the preceding proof, we can ask a question:
@@
Similar question can be asked for filled-in Julia sets as well, for, it was shown in Lemma 2.3 that $K(f\circ g)\subset K(f)\cap K(g)$. And the answer is: Let $w_{0}\in K(f\circ g)\ \backslash\ K(f)\cap K(g)$. These points won’t intersect with $I(f\circ g)$ and neither with $BU(f\circ g)$ as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset.$ Similarly, for intersection of filled-in Julia sets with escaping sets. Hence, we conclude no such points are possible. Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
So given the following results:
@@
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values
@@ 3. Some results for functions of class $\mathcal{B}$
Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$. For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11]. However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10]. We also know that $g(J(f))=J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
@@
$g(BU(f))\subset BU(f)\cup RP(f)$ for functions in class $\mathcal{B}.$
@@
If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded. Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions). Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below.
@@
$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}\text{ which also means }f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}\text{ which also means }f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
@@
$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}\text{ which also means }f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$. As $z_{0}\in J(f)$ as well and we know that $g(J(f))=J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case. Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$. Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=g(z_{0})$. So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
Case 2. When at least one of the subsequence still remains unbounded. Let that be the $k-$th subsequence i.e, $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e. $f^{n_{k}}(g(z_{0}))\rightarrow\infty$. $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e. $f^{m}(g(z_{0}))\rightarrow l_{1}$. This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$. Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$. ∎
@@
$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of class $\mathcal{B}$
@@
First, we assume that $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded. As can be seen, there are no subsequences that are unbounded. Similarly, there are no such points in $K(f)$. Therefore, $z_{0}\not\in K(f)\cup K(g)$.
@@
(3.1) $|f(g^{n})(z_{0})|\rightarrow l\ \text{for some}\ l$
(3.2) $|f(g^{n})(z_{0})|\rightarrow\infty$
where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$. If we have (3.2), then from the array of sequences argument, we see that all subsequences tend to infinity. If we have (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$. Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded. The third case is also not possible. This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded. Therefore, $z_{0}\not\in I(g)$. On similar lines, $z_{0}\not\in I(f)$. Therefore, $z_{0}\not\in I(f)\cup I(g)$. Thus, we see that there are no points s.t, $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. ∎
@@
The motivation for the proof comes from [14]. Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$. By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$. Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ (3.5). When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction. If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it. In the end, it will either be an entirely bounded sequence or there would be (un)bounded subsequence. The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index). But by definition escaping sets must have all the subsequences tending to infinity. Hence, we arrive at a contradiction. On similar lines, it can be shown when $z_{0}\in BU(g)$ we again arrive at a contradiction. So, $z_{0}\not\in BU(f)\cup BU(g)$.
@@
$RP(f\circ g)\supset RP(f)\cup RP(g)$ for commuting functions of class $\mathcal{B}$.

### v3

@@ 1. Introduction
When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
@@
There is yet another classification of these points based on the nature of their orbits; viz., the filled- in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. $K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]). The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain. Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions. We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is again transcendental.
@@
Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$. The following results are proven in this paper:
@@
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed functions is contained in the union of the bungee set of functions used for composition. The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the intersection of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cap RP(g)$.
@@
We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18]. Also, $\mathcal{S}\subset\mathcal{B}$. $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, in this case, the result clearly holds.
@@ 1.1. Observations
    Both sets of composite function is contained in union of individual functions (to be proved later on).
@@ 2. Main results
To begin with the commuting entire functions, we first understand the nature of these sets when functions have no finite asymptotic values in their singular set.
@@
Suppose $f$ and $g$ are commuting entire functions having no finite asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
@@
$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, that proves the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values. Similarly $K(g)$ is also completely invariant under $f$ as well.
Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument $K(f)$ is completely invariant under $g$ and from Theorem 2.1 $I(f)$ is completely invariant under $g.$ Hence, $BU(f)$ is also completely invariant under $g$. Similarly, $BU(g)$ will be completely invariant under $f$. ∎
@@
The filled-in Julia set of the composite function $f\circ g$ cannot be larger than the intersection of filled-in Julia set of the individual functions which are commuting and having no finite asymptotic values.
@@
We assume that $K(f\circ g)\supset K(f)\cap K(g)$. Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$. If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, then iterate of $z_{0}$ under $(f\circ g)^{n}$ cannot be bounded because $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory or will tend to $\infty$ (Theorem 2.2 and Theorem 2.1).
Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$. Now, as $w_{0}\not\in K(g)$, and after interchanging roles of $f$ and $g$ due to commutation, we observe that $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded (as $w_{0}$ will either be in $BU(g)$ or $I(g)$ and $f(BU(g)\subset BU(g)$, $f(I(g)\subset I(g)$). Hence, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})=f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ in which RHS is bounded while LHS is not which is a contradiction. Similar arguments may be given when $w_{1}\in B$. Hence, $K(f\circ g)\not\supset K(f)\cap K(g)$. ∎
The next result provides condition for equality of bungee set, filled-in Julia set and escaping set for commuting functions.
@@
We first establish that $BU(f\circ g)\subset BU(f)\cap BU(g).$ Let $z_{0}\in BU(f\circ g)$. By definition, there are two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ s.t.,
$|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R\ \text{for some}\ R$
$|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty$
So, for $\{m_{k}\}$ subsequence, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$. We observe same thing for $f$ as well i.e., $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$ (because $|f^{m_{k}}\circ g^{m_{k}}(z_{0})|=|g^{m_{k}}\circ f^{m_{k}}(z_{0})|<R$ as $f,g$ are commuting).
Now for $\{n_{k}\}$ subsequence, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ itself exhibits bungee behaviour. If we have the second case, it means $z_{0}\in K(g)$ (as this will imply all subsequences of $g^{j}(z_{0})$ are bounded where $j$ is some arbitrary index unrelated to $m_{k}$ and $n_{k}$). As a result, $g^{j}\circ f^{2j}\circ g^{j}(z_{0})$ will remain bounded (because $g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded). Hence, this case isn’t possible and $z_{0}\not\in K(g)$. On similar lines, it can be seen that $z_{0}\not\in K(f)$ where similar arguments follow after interchanging roles of $f$ and $g$ due to commutation. Now, when we have either of the first or the third case, it means the same thing. Both these cases essentially implies that $z_{0}\in BU(g)$. From Theorem 2.2, we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected. Due to commutation, we also conclude that $z_{0}\in BU(f)$ as well. Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
Given the preceding proof, we ask a question:
@@
Similar question can be asked for filled-in Julia sets as well, (as it was shown in Lemma (2.3) that $K(f\circ g)\subset K(f)\cap K(g)$). And the answer is: Let $w_{0}\in K(f)\cup K(g)\ \backslash\ K(f)\cap K(g)$. These points won’t intersect with $I(f\circ g)$ (as $I(f\circ g)\subset I(f)\cup I(g)$ and $K(f)\cup K(g)\cap I(f)\cup I(g)=\emptyset$ [14]) nor with $BU(f\circ g)$ (as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset)$ So, they will intersect with $K(f\circ g)$ which means $w_{0}\in K(f)\cap K(g)$ which is a contradiction. Hence, we conclude no such points are possible. Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
Hence, given the following results:
@@
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values.
@@ 3. Some results for functions of class $\mathcal{B}$
Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$. For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11]. However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10]. We also know that $g(J(f))\subset J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
@@
$g(BU(f))\subset BU(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
@@
If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded. Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions). Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below:
@@
$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}=f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}=f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
@@
$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}=f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$. As $z_{0}\in J(f)$ and we know that $g(J(f))\subset J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case. Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$. Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=z_{0}$. So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
Case 2. When at least one of the subsequence still remains unbounded. Let that be the $k-$th subsequence i.e., $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e., $f^{n_{k}}(g(z_{0}))\rightarrow\infty$. $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e. $f^{m}(g(z_{0}))\rightarrow l_{1}$. This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$. Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$. ∎
@@
$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
@@
First, we assume that $z_{0}\in K(g)$ which in turn implies $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded. As can be seen, there are no subsequences that are unbounded. Similarly, there are no such points in $K(f)$. Therefore, $z_{0}\not\in K(f)\cup K(g)$.
@@
(3.1) $|f(g^{n}(z_{0}))|\rightarrow l\ \text{for some}\ l$
(3.2) $|f(g^{n}(z_{0}))|\rightarrow\infty$
where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$. If we have case (3.2), then from the array of sequences argument, we see that all subsequences of $(f\circ g)^{n}(z_{0})$ tend to infinity. If we have case (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$. Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded. The third case is also not possible. This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded. Therefore, $z_{0}\not\in I(g)$. On similar lines, $z_{0}\not\in I(f)$. Therefore, $z_{0}\not\in I(f)\cup I(g)$. Thus, we see that there are no points s.t., $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. ∎
@@
The motivation for the proof comes from [14]. Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$. By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$. Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ using Lemma 3.4. When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e., $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction. If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it. In the end, it will either be an entirely bounded sequence or there would be an (un)bounded subsequence. The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index). But by definition, escaping sets must have all the subsequences tending to infinity. Hence, we arrive at a contradiction. On similar lines, it can be shown that when $z_{0}\in BU(g)$ we again arrive at a contradiction. So, $z_{0}\not\in BU(f)\cup BU(g)$.
@@
$RP(f\circ g)\supset RP(f)\cap RP(g)$ for commuting functions of class $\mathcal{B}$.